# Structure of Classifier Boundaries: Case Study for a Naïve Bayes Classifier

### v2

@@ 5 Discussion
There are two themes in this paper: the nature of the boundary and the surrogate measures of uncertainty based on the boundary, especially Neighbor Similarity. Concerning the former, the takeaway message is “beware the boundary.” For our Bayes classifier, all of our analyses suggest that the boundary may comprise 20% or more of the input space. The boundary is both not sparse and geometrically complicated.
5.1 General Issues
@@
The extent to which lying on or near the boundary $\mathcal{B}$ can or should affect either classifier output or decisions based on it is clearly contextual. The first issue is whether to check boundary status for a given query. So long as the input space has graph structure, the issue is computational, not conceptual: simply check all the neighbors, which, as noted, is completely parallelizable. If one cares only about whether the query is on the boundary, the process can be terminated once a neighbor with a different result is found. However, given the size of the boundary, this strategy is almost surely too blunt. In many cases, both scientific and broader impacts merit a quantified measure such as Neighbor Similarity. Although in our read classification case study, the consequences of misclassification are not major, for classifiers such as SeqScreen (Balaji et al.,, 2022), which is designed to identify pathogenic sequences, the consequences of incorrect decisions may be dramatic, and the risks of accepting high-uncertainty classifier results without further investigation may be unacceptable. Not knowing the level of uncertainty is equally problematic.
@@
Furthermore, in the face of rampart concern about data quality in DNA databases (Commichaux et al.,, 2021; Langdon,, 2014; Steinegger and Salzberg,, 2020), and given that changing a data point to a neighbor, as done in (Karr et al.,, 2022) to measure data quality, is seemingly the most innocuous data quality problem, [^9] As least with respect to the accuracy dimension of data quality, which is but one of many dimensions (Karr et al.,, 2006). it seems prudent to care about neighbors.
@@
In Section 4, we showed that that Neighbor Similarity relates to both inherent measures of uncertainty for our Bayes classifier. From an operational perspective, Figure 12 is more important. It contains receiver operating characteristic (ROC) curves for two decision rules, one based on accepting decisions for which MAP exceeds a threshold, which is encoded by color, the other for the decision rule based on accepting decisions for which NS exceeds a threshold. The two ROC curves are essentially identical. For the Bayes classifier, NS is as effective as MAP for making principled decisions.
@@
Figure 12: Left: ROC curve for the decision rule based on thresholded values of MAP. Right: ROC curve for the decision rule based on thresholded value of NS.
5.2 Scientific Generalizability
There are, of course, questions of scientific generalizability: are the properties observed somehow specific to our read dataset, or descriptive of the classifier $C$ more generally? And do they apply to other classifiers?
5.2.1 Other Datasets
To address the first question, we ran the classifier on five additional datasets:
ReadsNew:
A dataset containing 2000 Adeno reads, 2000 COVID reads and 2000 SARS reads, all of length 101, generated by the Mason_simulator, which can be viewed as addressing the question of replicate variability. There is no reason to expect the classifier to perform differently on it than on the main dataset.
ReadsEcoli:
A dataset containing 6000 Mason_simulator-generated reads of length 101 from an E. coli genome.
Random6K:
[…]
ReadsMixed1:
A dataset combining the original dataset and 2000 elements of ReadsEcoli, which represents a situation that the classifier is not prepared for: the presence of a fourth source of reads.
ReadsMixed2:
A dataset combining the original dataset and 100 elements of ReadsEcoli, which represents an alternative situation that the classifier is not prepared for: contamination from a fourth source of reads. [^10] Contamination of DNA samples is a pervasive problem in multiple contexts, especially analysis of ancient DNA (Reich,, 2018).
The results, including the core dataset ReadsOriginal for comparison, appear in Table 6 and Figure 13. In Table 6, the three columns under “Decisions” are percentages of the classifier output, as are the two columns under “Boundary.” The first two rows, representing two sets of reads from Adeno, COVID and SARS with approximately equal distributions of the read source, do not differ, meaning the that classifier seems not to be vulnerable to replicate variability. The last two rows differ the most from ReadsOriginal, and both the Decision distribution and the Boundary distribution differ dramatically from the first two rows. For reasons that are not yet clear, ReadsEcoli and Random6K are overwhelmingly classified as Adeno. The hypothesis that this constitutes a “vote splitting” artifact of there being one Adenovirus genome and two Coronavirus genomes is demonstrably false. In a set of 30,000 random sequences, if classifier choices are restricted to Adeno and Corona, only 77 of the 27,492 sequences classified as Adeno change to the latter. But ReadsEcoli and Random6K also differ: using Hellinger distances and empirical $p$-values, the differences between the Decision distributions and Boundary distributions are massively significant.
The third and fourth rows in Table 6 are, as expected, midway between the first two rows and the last two rows. Note that, for ease of interpretation, row 3 contains ReadsMixed2—the case with 100 contaminating E. coli reads, which should be, and is, closer to ReadsOriginal and ReadsNew than ReadsMixed1, with 2000 E. coli reads.
Figure 13 contains the ECDFs of Neighbor Similarity for all six datasets. That for ReadsOriginal (blue) is essentially completely overplotted by those for ReadsNew (cyan) and ReadsMixed2 (green). Interestingly, but consistent with the Boundary columns in Table 6, ReadsEcoli is now the most different from ReadsOriginal, ReadsNew and ReadsMixed2, even though it contains “real” reads. We performed one- and two-sided Kolmogorov-Smirnov tests on selected pairs of the ECDFs in Figure 13. No difference involving ReadsOriginal, ReadsNew and ReadsMixed2 was remotely significant. On the other hand, differences between ReadsOriginal and all of ReadsMixed1, Random6K, and ReadsEcoli were massively significantly, in the latter two cases with $p$-values of zero.
Therefore, Boundary Status and Neighbor Similarity are usable as diagnostics that indicate when a classifier is being used on data that differ from what was expected. For the Bayes classifier, “what is expected” is incorporated in the prior and the likelihood functions (Karr et al.,, 2023). For many other classifiers, the training data represent “what is expected.” The question is being investigated.
Table 8: Summary of results for the Bayes classifier applied to alternative datasets. Columns for Decision and Boundary are percentages.
Decision Boundary
Dataset Count Adeno COVID SARS No Yes
ReadsOriginal 5869 32.94 34.09 32.97 69.48 30.52
ReadsNew 6000 32.43 33.82 33.75 68.7 31.3
ReadsMixed2 5969 33.86 33.59 32.55 69.69 30.31
ReadsMixed1 7869 48.1 26.2 25.7 74.1 25.9
ReadsEcoli 6000 92.767 2.933 4.300 87.35 12.65
Random6K 6000 90.8333 0.3833 8.7833 78.75 21.25
Figure 13: Empirical cumulative distribution functions of Neighbor Similarity for six datasets.
5.2.2 Other Classifiers
Full results will be reported in forthcoming papers, but initial experiments show similar behavior for other classifiers applied to the same dataset, including partition models, random forests and neural networks.
6 Conclusions
@@
The two principal contributions of this paper are surrogate measures of uncertainty for classifiers and methods for investigating their boundaries. In particular, we propose Neighbor Similarity as a measure of uncertainty for classifiers lacking inherent measures of uncertainty. Despite the computational burden, the potential to avoid incorrect decisions is enormous.
@@
As always, at least as many questions are raised as answered. One of the more challenging is whether Neighbor Similarity is viable in the context of production-grade classifiers such as SeqScreen. A central issue may be whether the information gained justifies the computational cost. A (still short) DNA sequence of length 10,000 has more than 40,000 neighbors, and running SeqScreen on all of them may be prohibitive. We showed the potential of sampling to estimate Neighbor Similarity from a subset of neighbors; investigating demonstrably effective strategies to do so is a priority.
@@
Underlying everything is a scientific and statistical issue not addressed in this paper and clearly needing more attention. Is a low value of $\mbox{NS}(R)$ a statement about the read or about the classifier? Or, if present, the training data? As long there is a range of observed values of NS, and especially if there are values near 1, then low values are plausibly statements about reads. On the other hand, uniformly small values of $\mbox{NS}(R)$ seem to suggest problems with the classifier (or, if present, the training data). But then, what if the problem is just innately hard?
@@ Appendix A Triplet Distributions of the Three Genomes
Appendix B Variable Importance for a Pruned Partition Model
Variable importance for a pertition model measures the extent to which each predictor affects the splits in the model. The values below are for the 182-terminal node model in Section 4.4.
Table 10: Relative importance of the 64 predictors for the “halve the error rate” pruning of the partition model in Section 4.4.
Variable Importance Variable Importance
TTT 4.5974 GGA 1.3454
TAT 4.5730 GAA 1.3437
GTT 3.6585 CCG 1.2929
GCG 3.6076 ACG 1.2341
AAG 2.7438 AAC 1.1639
CGG 2.6345 AGA 1.1582
CAC 2.5845 GAC 1.1186
ATA 2.5252 CAG 1.1119
TTA 2.4875 ATC 1.0779
CGC 2.4463 TCT 1.0733
CCC 2.1925 TAC 1.0558
AAA 2.0777 GCC 1.0429
TTG 2.0324 GAG 1.0367
GAT 2.0316 TGA 1.0167
ATT 2.0179 CAT 1.0009
ATG 1.8936 TCG 0.9881
TGT 1.8814 GTA 0.9734
AGG 1.7983 CGA 0.9728
TAA 1.7764 CCT 0.8902
ACA 1.7416 GTC 0.8836
GGC 1.7339 CTA 0.8758
AAT 1.7241 TGC 0.8617
CTT 1.7121 TCA 0.8203
TTC 1.6527 GGT 0.7777
GCA 1.6409 CTG 0.7758
AGC 1.6359 ACT 0.6747
CCA 1.5917 GCT 0.6221
AGT 1.5696 TGG 0.6044
ACC 1.5281 CTC 0.5921
CAA 1.4947 TCC 0.5856
TAG 1.3968 CGT 0.3860
GTG 1.3584 GGG 0.3021

### v3

@@ 5 Discussion
The extent to which lying on or near the boundary $\mathcal{B}$ can or should affect either classifier output or decisions based on it is contextual. The first issue is whether to check Boundary Status for a given query. If one cares only about whether the query lies on the boundary, the process can be terminated once a neighbor with a different result is found. However, given the size of the boundary, this strategy is almost surely too blunt. In many cases, both scientific and broader impacts merit a quantified measure such as Neighbor Similarity. Although in our read classification case study, the consequences of misclassification are arguably not major, for classifiers such as SeqScreen (Balaji et al.,, 2022), which is designed to identify pathogenic sequences, the consequences of incorrect decisions may be dramatic, and the risks of accepting high-uncertainty classifier results without further investigation may be too high in contexts. Not knowing the level of uncertainty is equally problematic.
@@
Furthermore, in the face of rampant concern about data quality in DNA databases (Commichaux et al.,, 2021; Langdon,, 2014; Steinegger and Salzberg,, 2020), and given that changing a data point to a neighbor, as done in (Karr et al.,, 2022) to measure data quality, is seemingly the most innocuous data quality problem (at least with respect to the accuracy dimension of data quality, which is but one of many dimensions (Karr et al.,, 2006)), it seems prudent to care about neighbors. This issue is discussed at length in Karr and Ruane, (2026).
@@
In Section 4, we showed that that Neighbor Similarity relates to both inherent measures of uncertainty for our Bayes classifier. From an operational perspective, Figure 11 is equally important. It contains receiver operating characteristic (ROC) curves for two decision rules, one based on accepting decisions for which MAP exceeds a threshold, which is encoded by color, the other for the decision rule based on accepting decisions for which NS exceeds a threshold. The two ROC curves are essentially identical. For the Bayes classifier, NS is as effective as MAP for making principled decisions. Although this point is addressed elsewhere (Porter and Karr,, 2024; Karr and Ruane,, 2026), we emphasize that NS is implementable for all classifiers.
@@
Figure 11: Left: ROC curve for the decision rule based on thresholded values of MAP. Right: ROC curve for the decision rule based on thresholded value of NS. In each panel, the $x$-axis is the False Positive Rate (1 - Specificity) and the $y$-axis is the True Positive Rate (Sensitivity).
6 Conclusion
@@
The three principal contributions of this paper are the concept of the boundary, surrogate measures of uncertainty for classifiers and methods for investigating boundaries. In particular, we propose Neighbor Similarity as a measure of uncertainty for classifiers lacking inherent measures of uncertainty. Despite the computational burden, the potential to avoid incorrect decisions is significant.
@@
As always, at least as many questions are raised as answered. One is whether Neighbor Similarity is viable in the context of production-grade classifiers such as SeqScreen. A central issue may be whether the information gained justifies the computational cost. A (still short) DNA sequence of length 10,000 has more than 40,000 neighbors, and running SeqScreen on all of them may be prohibitive. We showed the potential of sampling to estimate Neighbor Similarity from a subset of neighbors; investigating demonstrably effective strategies to do so is a priority.
@@
Underlying everything is a scientific and statistical issue not addressed in this paper and clearly needing more attention. Is a low value of $\mbox{NS}(R)$ a statement about the read or about the classifier? Or, if present, about the training data? As long there is a range of observed values of NS, and especially if there are values near 1, then low values are plausibly statements about reads. Stronger evidence that a read is problematic would be that it is on the boundary for multiple classifiers, which initial evidence shows to be true (Porter and Karr,, 2024). On the other hand, uniformly small values of NS seem to suggest problems with the classifier (or, if present, the training data). But then, what if the problem is just innately hard?